OpenStudy (ttp):

A waiter earned $16 in tips in an hour. He was paid in $1 bills and $5 bills. Use a model to determine how many $5 bills he received if he received a total of 12 bills. 1 2 3 4

OpenStudy (anonymous):

4

OpenStudy (anonymous):

$16=$1n+$5m 12=n+m rearrange for n n=12-m sub back value into:$16=$1n+$5m $16=$1(12-m)+$5m $16=$12-$m+$5m 4m=4 m=1 therefore 1 five dollar bill and 11 1 dollar bills